12. Looking at the increasing trend of smoking habit of teenagers, Ministry of Health, research team undertake a study and found that the price elasticity of demand for cigarettes is about 0.75. If a pack of cigarettes currently costs Rs.75 and government wants to reduce smoking by 30 percent, by how much should it increase the price?


1
Expert's answer
2022-02-28T09:38:57-0500

If price elasticity of ci garettes is at 0.75 and the government want to reduce consumption by:

"0.75\times increase = 30;"

"Increase=\frac{30}{0.75}=40"

Increasing price by 40% will reduce consumption by 30%.

"New price=1.4\times75=105"

Increase=105-75=Rs. 30


For consumption to be reduced to 30%, price should be increased by Rs. 30 which is 40% increase.